# DiffSSC: Semantic LiDAR Scan Completion using Denoising Diffusion Probabilistic Models

### v2

@@ V Conclusions and Outlook
We proposed DiffSSC, a novel SSC approach based on a diffusion model. It takes raw LiDAR point clouds as input and jointly predicts missing points along with their semantic labels, thereby extending the application boundaries of diffusion models. We evaluated our method on two autonomous driving datasets, achieving performance that surpasses the state-of-the-art. In future work, we will explore methods to enhance inference speed by streamlining the step-by-step inference process, enabling the application of diffusion models  [35]. Regarding the impact of noise schedules, we will also explore more complex yet efficient scheduling mechanisms, such as adaptive schedules [36].

### v3

@@ V Conclusions and Outlook
We proposed DiffSSC, a novel SSC approach based on DDPMs. It takes raw LiDAR point clouds as input and jointly predicts missing points along with their semantic labels, thereby extending the application boundaries of diffusion models. We evaluated our method on two autonomous driving datasets, achieving state-of-the-art performance. In future work, we plan to explore integrating cross-modal signals and prompt-guided learning to enhance scene understanding [36, 37, 38]. We will also explore strategies for improving inference efficiency and reducing resource consumption, inspired by recent advances in lightweight modeling [39, 40].